20grams of NaOH is dissolved in enough water to make 1 liter of solution
1
Expert's answer
2020-11-11T10:32:25-0500

m (NaOH) = 20 g

n (NaOH) = m (NaOH)/M (NaOH) = 20 / (23 + 16 + 1) = 20 / 40 = 0.5 mol

C (NaOH) = n (NaOH) / V(solution) = 0.5 mol / 1 L = 0.5 mol/L